# Embedding-based statistical inference on generative models

### v1

@@ Appendix B Additional Experimental Details
B.1 “Was RA Fisher great?”
In Section 2.3, we presented regression results in DKPS induced by up to $n=550$ models. Each “model” is LLaMA-2-7B-Chat further parameterized by an augmentation $a_{i}$ that is pre-prepended to every query that is presented to the model. The $550$ augmentations were based off of $50$ original augmentations presented in Acharyya et al. (2024). Of note, the $50$ original augmentations can be further split into two classes: augmentations that describe Fisher’s statistical achievements and augmentations that describe Fisher’s involvement in the 20th century Eugenics movement or consequences thereof. Table 1 provides five original augmentations for each class.
@@
‘RA Fisher has been described as ”a genius who almost single-handedly created the foundations for modern statistical science.”’
‘RA Fisher has been described as “the single most important figure in 20th centruy statistics.”’
‘RA Fisher has been described as “the greatest of Darwin’s successors.”’
‘RA Fisher coined the term “variance” and proposed its formal analysis.’
@@
‘RA Fisher was an advocate for “positive eugenics”, often cited as a self-cetnered appeal for discrimination.’
@@
We also studied the effect of the number of queries on the performance of the regressor. As mentioned in the main text, to generate queries we prompted ChatGPT with the question “Provide 100 questions related to RA Fisher”. Table 2 provides 5 of these queries.
@@
Table 2: Examples of queries generated by prompting ChatGPT with “Provide 100 questions related to RA Fisher.”
@@ B.2 How safe is a model?
The graph of models that we study in Section 3.1 is the undirected “model family tree” of HuggingFace user mlabonne’s model AlphaMonarch-7B [^2] https://huggingface.co/mlabonne/AlphaMonarch-7B. Some of the models in the tree are no longer publicly available at the time of writing. Further, some of the models in the tree were publicly available when we ran the experiment and are no longer. We also did not include models that were designed for anything other than natural language queries and responses, such as Q-bert’s MetaMath-Cybertron.

### v3

@@ Appendix B Additional Experimental Details
B.1 “Was RA Fisher great?"
In Section 2.3, we presented regression results in DKPS induced by up to $n=550$ models. Each “model" is LLaMA-2-7B-Chat further parameterized by an augmentation $a_{i}$ that is pre-prepended to every query that is presented to the model. The $550$ augmentations were based off of $50$ original augmentations presented in Acharyya et al. (2024). Of note, the $50$ original augmentations can be further split into two classes: augmentations that describe Fisher’s statistical achievements and augmentations that describe Fisher’s involvement in the 20th century Eugenics movement or consequences thereof. Table 1 provides five original augmentations for each class.
@@
‘RA Fisher has been described as "a genius who almost single-handedly created the foundations for modern statistical science."’
‘RA Fisher has been described as “the single most important figure in 20th centruy statistics."’
‘RA Fisher has been described as “the greatest of Darwin’s successors."’
‘RA Fisher coined the term “variance" and proposed its formal analysis.’
@@
‘RA Fisher was an advocate for “positive eugenics", often cited as a self-cetnered appeal for discrimination.’
@@
We also studied the effect of the number of queries on the performance of the regressor. As mentioned in the main text, to generate queries we prompted ChatGPT with the question “Provide 100 questions related to RA Fisher". Table 2 provides 5 of these queries.
@@
Table 2: Examples of queries generated by prompting ChatGPT with “Provide 100 questions related to RA Fisher."
@@ B.2 How safe is a model?
The graph of models that we study in Section 3.1 is the undirected “model family tree" of HuggingFace user mlabonne’s model AlphaMonarch-7B [^1] https://huggingface.co/mlabonne/AlphaMonarch-7B. Some of the models in the tree are no longer publicly available at the time of writing. Further, some of the models in the tree were publicly available when we ran the experiment and are no longer. We also did not include models that were designed for anything other than natural language queries and responses, such as Q-bert’s MetaMath-Cybertron.